OpenStudy (anonymous):

how would i solve this equation?? 4(x - 6) = -2(x - 3)

OpenStudy (aroub):

You first distribute, 4x-24=-2x+6 then put the like terms on one side and the other like terms on the other side, 4x+2x=24+6 now just add them, 6x=30 now multiply 1/6 by both sides to get rid of the 6, x= 5

OpenStudy (anonymous):

exactly

OpenStudy (aroub):

:D

OpenStudy (anonymous):

Could you help me with my Elasticity question?